OpenStudy (anonymous):

Solve: 4(x + 6) = 6(x − 2)

OpenStudy (anonymous):

4x+24 = 6x-12 6x-4x = 24+12 2x = 36 x =?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No probs!